OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (–10, –5); m = 1/6 a. y + 10 = (1/6)(x + 5) b. y + 5 = (1/6)(x + 10) c. y – 5 = (1/6)(x – 10) d. y – 5 = (1/6)(x + 10)

OpenStudy (anonymous):

point slope form is y - yo = m ( x - xo) plug in y-(-5) = 1/6 ( x-(-10)) y+5 = (1/6)(x+10) B that is.